How many liters each of a 40​% antifreeze solution and a 80​% antifreeze solution must be mixed to make 8 liters of a 50​% antifreeze​ solution?
Answer by VFBundy(438)

Liters of 40% solution = x
Liters of 80% solution = y

x + y = 8
0.40x + 0.80y = 0.50(8)

x + y = 8
0.40x + 0.80y = 4

Multiply first equation by -0.80:

-0.80 - 0.80y = -6.40
0.40x + 0.80y = 4

Add the equations together:

-0.40x = -2.40

x = 6

Since x = 6, and we know x + y = 8, that means y = 2.

Liters of 40% solution = x = 6
Liters of 80% solution = y = 2